ok here it is it is a mixture problem. the formula is V+AC value amount cost.
A clothing company buys 60 yards of cotton and 90 yards of wool fro a total cost of $1800. Another purchase at the same price included 80 yards of cotton and 20 yds of wook for a total cost of $1000. Find the cost per yard of the cotton and wool?

Answer by stanbon(57984)

A clothing company buys 60 yards of cotton and 90 yards of wool fro a total cost of $1800. Another purchase at the same price included 80 yards of cotton and 20 yds of wook for a total cost of $1000. Find the cost per yard of the cotton and wool?
---------------------------------------
Value Equation: 60c + 90W = 1800 dollars
Value Equation: 80c + 20w = 1000 dollars
---------------------
6c + 9w = 180
8c + 2w = 100
------------------
Modify:
12c + 18w = 360
72c + 18w = 900
---------------------
60w = 540
w = $9 (price for a yard of wool)
---
Substitute into 6c + 9w = 180 to solve for "c":
6c + 9*9 = 180
6c = 99
c = $16.50 (price for a yard of cotton
=============================
Cheers,
Stan H.